OpenStudy (anonymous):

#9 find the function value f(x)=x^3 f(7)=

OpenStudy (anonymous):

7^3=343

OpenStudy (anonymous):

7^3=7*7*7=343

OpenStudy (anonymous):

nancy is right on that one

OpenStudy (anonymous):

lol im dumb